OpenStudy (anonymous):

three-fifths of a number decreased by one is twenty-three.what is the number?

OpenStudy (amistre64):

n(3/5) - 1 = 23

OpenStudy (anonymous):

x*(3/5) - 1 = 23 x*(3/5) = 24 x = 24*5/3 x = 40

OpenStudy (anonymous):

i think its actually 3/5(n-1)

OpenStudy (anonymous):

which is 118/3 but if they are right then my english skills are just off

OpenStudy (amistre64):

english to math translation is hard; but your interp is plausible